# X-Dancer: Expressive Music to Human Dance Video Generation

### v1

@@ Abstract
We present X-Dancer, a novel zero-shot music-driven image animation pipeline that creates diverse and long-range lifelike human dance videos from a single static image. As its core, we introduce a unified transformer-diffusion framework, featuring an autoregressive transformer model that synthesize extended and music-synchronized token sequences for 2D body, head and hands poses, which then guide a diffusion model to produce coherent and realistic dance video frames. Unlike traditional methods that primarily generate human motion in 3D, X-Dancer addresses data limitations and enhances scalability by modeling a wide spectrum of 2D dance motions, capturing their nuanced alignment with musical beats through readily available monocular videos. To achieve this, we first build a spatially compositional token representation from 2D human pose labels associated with keypoint confidences, encoding both large articulated body movements (e.g., upper and lower body) and fine-grained motions (e.g., head and hands). We then design a music-to-motion transformer model that autoregressively generates music-aligned dance pose token sequences, incorporating global attention to both musical style and prior motion context. Finally we leverage a diffusion backbone to animate the reference image with these synthesized pose tokens through AdaIN, forming a fully differentiable end-to-end framework. Experimental results demonstrate that X-Dancer is able to produce both diverse and characterized dance videos, substantially outperforming state-of-the-art methods in term of diversity, expressiveness and realism.
@@
Figure 1: We present X-Dancer, a unified transformer-diffusion framework for zero-shot, music-driven human image animation from a single static image, capable of handling diverse body forms and appearances. Our method enables the synthesis of highly expressive and diverse full-body dance motions that are synchronized with music, including detailed movements at the head and hands, which are then seamlessly translated into vivid and lifelike dance videos. Code and model will be available for research purposes.
@@ 1 Introduction
With recent advances in diffusion models that have achieved unprecedented diversity and quality in image and video generation, numerous works have leveraged their generative capabilities to synthesize visually compelling videos by animating a reference image with motion signals such as 2D skeletons [15, 6], dense poses [43], and depth maps [52]. Unlike motion transfer settings that derive motion from a driving video, our goal is to generate motion signals that are consistent with the reference body shape and aligned with the input musical beats. Recently, a few studies [42, 35] have attempted to synthesize visual outputs end-to-end from audio inputs using diffusion models. While these methods have advanced in realism and dynamic quality, they still struggle to capture long-range motion and audio context due to high computational demands. Moreover, these frameworks have primarily been validated on audio-driven portrait animations, leaving uncertain whether they can accommodate the complexities of full-body kinematics and rapid motion transitions.
@@
To this end, we propose X-Dancer, a unified framework that integrates an autoregressive transformer model for generating extended dance sequences attuned to input music, coupled with a diffusion model to produce high-resolution, lifelike videos. In contrast to prior methods focused on 3D music-to-dance motion generation, our approach models 2D human motion, leveraging widely accessible dance videos where 2D pose estimation is both reliable and scalable. For effective autoregressive motion generation, we develop a multi-part tokenization scheme for per-frame 2D whole-body poses, incorporating detected keypoint confidences to capture multi-scale human motions with motion blur and various occlusions. Thereafter we train a cross-modal transformer that auto-regressively predicts future N-frame pose tokens, paired with per-frame music features extracted with Jukebox [9] and Librosa [19]. Our design enables the model to capture a broader diversity of expressive, music-aligned movements with enhanced scalability in both model complexity and data scale. We then leverage a T2I diffusion model with temporal layers to animate the reference image by implicitly translating the generated confidence-aware motion tokens into spatial guidance via a trainable motion token decoder. Specifically, it upsamples from a learnable feature map into multi-scale spatial feature guidance, integrating motion tokens through AdaIN [16] during the upsampling process. By co-learning pose translation with temporal and appearance reference modules [12, 5], the diffusion backbone interprets motion tokens with temporal context and body shape reference, leading to better shape-disentangled pose control, smoother and more robust visual outputs even under low-confidence or jittering pose sequences. This design establishes an end-to-end transformer-diffusion learning framework, merging the transformer’s strengths in long-context understanding and generation with the diffusion model’s high-quality visual synthesis.
@@
To the best of our knowledge, X-Dancer represents the first music-driven human image animation framework. Trained on a curated dataset of 100K music-dance video clips, our method excels at generating diverse, expressive and detailed whole-body dance video animations attuned to input music, adaptable to both realistic and stylized human images across various body types. We comprehensively evaluate our model on challenging benchmarks, and X-Dancer outperforms state-of-the-art baselines both quantitatively and qualitatively. Additionally, we highlight its scalability and customization capabilities, showcasing scalability across different model and data scales, as well as fine-tuning to specific choreography. We summarize our contributions as follows,
@@ 2.1 Music to Dance Generation
Significant progress has been made in realistic human motion generation [51, 32, 37, 27, 7, 48] from various inputs, such as input speech [11, 30, 25, 47, 2] or text [34, 17, 29, 50]. However, the task of music-to-dance generation [32, 1, 37, 24, 23, 20, 28, 46] presents unique challenges: (1) ensuring the generated dance rhythmically aligns with the music, and (2) producing intricate motions with diverse styles and speeds. Due to data limitations, several 3D human pose datasets [21, 22, 20, 28] have been proposed. The AIST++ dataset [21] is a 3D music pose dataset containing 1,408 dance motions tailored to various music genres. FineDance [22], a 3D dataset focusing on fine-grained hand motions, includes 14 hours of dance data. Models like Bailando [32] and EDGE [37] have leveraged AIST++ for training. Bailando [32] pioneered using a VQ-VAE for 3D pose encoding, followed by an Actor-Critic GPT model to generate body poses conditioned on music. EDGE [37] applied a diffusion framework to predict human poses from a noisy sequence, conditioning on music and employing long-form sampling for extended dance sequences. Despite these advancements, existing approaches trained on 3D datasets often generate dance poses with limited diversity and are primarily confined to a 3D format. X-Dancer overcomes these data constraints by leveraging a broad spectrum of 2D dance motions, aligning complex movements with musical beats, and utilizing widely available video data to enhance scalability.
@@ 2.2 Human Image Animation
With the advancements in diffusion models [14, 12, 4], generating high-quality human videos has become feasible. The introduction of ControlNet [49] and PoseGuider [15] has further empowered prior methods [15, 6, 43, 52] to create realistic dance videos from a reference image using motion signals such as 2D skeletons, dense poses, and depth maps. While these methods provide explicit control over body poses and allow flexible motion manipulation, they often struggle to preserve the original body shape and other identity-specific features, leading to noticeable identity leakage, particularly when significant differences exist between the target and source. Several methods [42, 18, 35] adopt an end-to-end pipeline for generating realistic talking human videos from audio by injecting audio features directly into the network via cross-attention layers. However, while these methods excel in capturing micro-expressions and lip movements, they fall short in handling dynamic, shape-preserving dance movements. X-Dancer uses a diffusion backbone with AdaIN to animate the reference image in an end-to-end framework, enabling high-quality, shape-consistent dance generation while effectively preserving the identity of the reference image.
Figure 2: Overview of X-Dancer. We propose a cross-conditional transformer model to autoregressively generate 2D human poses synchronized with input music, followed by a diffusion model that produces high-fidelity videos from a single reference image $I_{R}.$ First, we develop a multi-part compositional tokenization for 2D poses, encoding and quantizing each body part independently with keypoint confidence. These tokens are then merged into a whole-body, confidence-aware pose using a shared decoder. Next, we train a GPT-based transformer to autoregressively predict future pose tokens with causal attention, conditioned on past poses and aligned music embeddings. For global music style and motion context, we incorporate the entire music sequence and sampled prior poses. With a learnable motion decoder, we generate multi-scale spatial pose guidance upsampled from a learned feature map, incorporating the generated motion tokens within a temporal window (16 frames) using AdaIN. By co-training the motion decoder and temporal modules, our diffusion model is capable of synthesizing temporally smooth and high-fidelity video frames, while maintaining consistent appearance with the reference image with a trained reference net.
@@ 3.1 Data Representations
To generate human dance videos, we employ Latent Diffusion Models [31] that synthesize samples in the image latent space, facilitated by a pretrained image auto-encoder. During training, latent representations of images are progressively corrupted with Gaussian noise $\epsilon,$ following the Denoising Diffusion Probabilistic Model (DDPM) framework [13, 33]. A UNet-based denoising framework, enhanced with spatial and temporal attention blocks, is then trained to learn the reverse denoising process.
@@
For training, we apply human-centric cropping to the videos, yielding a unified resolution of $896\times 512,$ that encompasses half-body to full-body dances. Instead of directly capturing the intricate pixel-wise movements on music conditions, we first establish the correlation between music and human dance movement, which subsequently guides the final visual synthesis.
Motion Modeling.
@@
In contrast to prior approaches that generate human motions in 3D space, we represent diverse dance motions as 2D pose sequences. Compared to its 3D counterparts, 2D human dance motions are widely accessible from large collections of monocular videos, eliminating the need for complex multiview capture setups or labor-intensive 3D animations. Furthermore, the detection of 2D poses is significantly more reliable and robust. To enhance the realism and motion expressiveness, we model not only large body articulations but also finer details such as head movements and hand gestures, capturing subtle nuances in motion. Notably, we incorporate keypoint detection confidence into our pose representation, allowing the model to account for motion blur and occlusions. Each per-frame whole-body pose with $J$ joints is thus represented as $p\in\mathcal{R}^{J\times 3},$ where the last dimension encodes the keypoint confidence.
Music Embedding.
@@
Inspired from [37], we utilize the pre-trained Jukebox model [9] to extract rich musical features, supplemented by rhythmic information with one-hot encoding of music beats using an audio processing toolbox Librosa [19]. We resample and synchronize these extracted embeddings—denoted as $F^{J}_{1:T}$ for Jukebox and $F^{L}_{1:T}$ for Librosa—to the video frame rate, ensuring per-frame alignment between the music and visual elements.
3.2 Transformer-Based Music-to-Dance Motion Generation
@@
Given a collection of monocular, single-person, music-synchronized dance videos with paired 2D whole-body pose detections, we aim to model the intricate, non-linear correlation between skeletal dance motion and musical features. To achieve this, we first introduce a compositional, confidence-aware VQ-VAE, which captures diverse and nuanced human dance poses across different body parts. Next, we leverage a GPT-like transformer to autoregressively predict future pose tokens, modeling temporal motion transitions within the token space and aligning them with synchronized music embeddings.
@@
Compositional Confidence-Aware Human Pose Tokenization. Our approach builds on the standard VQ-VAE framework, trained in a self-supervised manner. Given whole-body 2D poses with associated keypoint confidences $p$, an 1D convolutional encoder maps $p$ into a latent pose embedding $z_{e}(p)=E(p),$ which is then quantized by mapping to its nearest representation $z_{q}(p)$ within a learnable codebook, and finally the decoder $D$ reconstructs the pose $\hat{p}$ from $z_{q}(p).$ However, as prior studies [32, 45, 41] suggest, the dependencies between spatial keypoints are complex, and a vanilla VQ-VAE often struggles to capture subtle pose details, such as finger movements and head tilts, due to information loss during quantization and multi-frequency nature of pose variations across different body parts.
@@
To improve expressive coverage, we train independent 2D pose encoders $E^{j}$ and learn $B=5$ separate codebooks $Z^{j}_{q}$ for upper and lower half bodies, left and right hands, and head respectively, allowing the model to spatially decompose 2D whole-body pose variations across different frequency levels. By partitioning poses this way, distinct body-part codes can be flexibly combined, enhancing the range of expressiveness represented within each individual codebook. To capture part-wise spatial correlations and ensure information flow across body parts, we concatenate the quantized pose latents and feed them into a shared decoder. The resulting embedding is then mapped to reconstructed keypoint coordinates through separate projection heads, enabling joint reconstruction while preserving nuanced part-specific details.
@@ 3.3 Diffusion-Based Dance Video Synthesis
With the generated motion token sequences from our trained transformer model, we employ a diffusion model to synthesize high-resolution, realistic human dance videos, conditioned on the generated motions and a given reference image. To achieve this, we leverage a pretrained T2I diffusion backbone [31] and incorporate additional temporal modules [12] to improve cross-frame consistency and ensure temporal smoothness. For transferring the reference image context, a reference network, as a trainable copy of the backbone UNet, extracts reference features of identity appearance and background which are cross-queried by the self-attentions within the backbone UNet. Motion control is achieved through an addition module, often configured in ControlNet [49] or light-weight PoseGuider [15], which translates the motion conditions into 2D spatial guidance additive to the UNet features.
@@
In place of explicit translation of motion tokens into 2D skeleton maps as pose guider conditions, we introduce a trainable motion decoder that implicitly and directly translates the 1D pose tokens into 2D spatial guidance, added towards the intermediate features of the denoising UNet. Starting from a learnable 2D feature map, the decoder injects the 1D pose token sequences including the keypoint confidences with AdaIN layers [16], progressively upsampling this feature map into multiple scales aligned with the resolutions of the denoising UNet features. The motion decoder is trained alongside the temporal module within a 16-frame sequence window, effectively decoding token sequences into continuous pose guidance that integrates temporal context from adjacent frames. Moreover, by incorporating reference image context during training, we observe empirically that the synthesized pose guidance retains minimal identity and body shape information compared to skeleton-based pose guiders, enabling generated pose tokens to adapt seamlessly to characters with varied body shapes and appearances.

### v2

@@ Abstract
We present X-Dancer, a novel zero-shot music-driven image animation pipeline that creates diverse and long-range lifelike human dance videos from a single static image. As its core, we introduce a unified transformer-diffusion framework, featuring an autoregressive transformer model that synthesize extended and music-synchronized token sequences for 2D body, head and hands poses, which then guide a diffusion model to produce coherent and realistic dance video frames. Unlike traditional methods that primarily generate human motion in 3D, X-Dancer addresses data limitations and enhances scalability by modeling a wide spectrum of 2D dance motions, capturing their nuanced alignment with musical beats through readily available monocular videos. To achieve this, we first build a spatially compositional token representation from 2D human pose labels associated with keypoint confidences, encoding both large articulated body movements (e.g., upper and lower body) and fine-grained motions (e.g., head and hands). We then design a music-to-motion transformer model that autoregressively generates music-aligned dance pose token sequences, incorporating global attention to both musical style and prior motion context. Finally we leverage a diffusion backbone to animate the reference image with these synthesized pose tokens through AdaIN, forming a fully differentiable end-to-end framework. Experimental results demonstrate that X-Dancer is able to produce both diverse and characterized dance videos, substantially outperforming state-of-the-art methods in term of diversity, expressiveness and realism. See our project page for more results: zeyuan-chen.com/X-Dancer/
@@
Figure 1: We present X-Dancer, a unified transformer-diffusion framework for zero-shot, music-driven human image animation from a single static image. Our approach accommodates diverse body forms and appearances, generating diverse and highly expressive synchronized full-body dance motions. It captures both fine-grained movements of the head and hands and large-scale actions such as body rotations and jumps, seamlessly translating them into vivid and lifelike dance videos.
@@ 1 Introduction
With recent advances in diffusion models, numerous works have leveraged their generative capabilities to synthesize visually compelling videos by animating a reference image with motion signals such as 2D skeletons [17, 7, 18], dense poses [54], and depth maps [66]. Unlike motion transfer settings that derive motion from a driving video, our goal is to generate motion signals that are consistent with the reference body shape and aligned with the input musical beats. Recently, a few studies [53, 42] have attempted to synthesize visual outputs end-to-end from audio inputs using diffusion models. While these methods have advanced in realism and dynamic quality, they still struggle to capture long-range motion and audio context due to high computational demands. Moreover, these frameworks have primarily been validated on audio-driven talking heads, leaving uncertain whether they can accommodate the complexities of full-body kinematics and rapid motion transitions.
@@
To this end, we propose X-Dancer, a unified framework that integrates an autoregressive transformer model for generating extended dance sequences attuned to input music, coupled with a diffusion model to produce high-resolution, lifelike videos. In contrast to prior methods focused on 3D music-to-dance motion generation, our approach models 2D human motion, leveraging widely accessible dance videos where 2D pose estimation is both reliable and scalable. For effective autoregressive motion generation, we develop a multi-part tokenization scheme for per-frame 2D whole-body poses, incorporating detected keypoint confidences to capture multi-scale human motions with motion blur and various occlusions. Thereafter we train a cross-modal transformer that auto-regressively predicts future N-frame pose tokens, paired with per-frame music features extracted with Jukebox [11] and Librosa [22]. Our design enables the model to capture a broader diversity of expressive, music-aligned movements with enhanced scalability in both model complexity and data scale. We then leverage a T2I diffusion model with temporal layers to animate the reference image by implicitly translating the generated confidence-aware motion tokens into spatial guidance via a trainable motion token decoder. Specifically, it integrates motion tokens through AdaIN [19] along upsampling from a learnable feature map into multi-scale spatial feature guidance. By co-learning pose translation with temporal and appearance reference modules [14, 6], the diffusion backbone interprets motion tokens with temporal context and body shape reference, leading to better shape-disentangled pose control, smoother and more robust visual outputs even under low-confidence or jittering pose sequences. This design establishes an end-to-end transformer-diffusion learning framework, merging the transformer’s strengths in long-context understanding and generation with the diffusion model’s capability in high-quality visual synthesis.
@@
To the best of our knowledge, X-Dancer represents the first music-driven human image animation framework. Trained on a large curated music-dance video dataset, our method excels at generating diverse, expressive and detailed whole-body dance videos attuned to input music, adaptable to both realistic and stylized human images across various body types. We extensively evaluate our model on challenging benchmarks, and X-Dancer outperforms state-of-the-art baselines both quantitatively and qualitatively. Additionally, we highlight its scalability and customization capabilities, showcasing scalability across varying model and data scales, as well as fine-tuning to characterized choreography. We summarize our contributions as follows,
@@ 2.1 Music to Dance Generation
Significant progress has been made in realistic human motion generation [65, 38, 44, 31, 9, 61] from various inputs, such as speech [13, 36, 29, 60, 2] or text [41, 20, 35, 63]. However, the task of music-to-dance generation [38, 1, 44, 28, 27, 24, 32, 59, 51, 67, 55] presents unique challenges: (1) ensuring the generated dance rhythmically aligns with the music, and (2) producing intricate motions with diverse styles and speeds. There has been several 3D human pose datasets [25, 26, 24, 32] proposed. The AIST++ dataset [25] is a 3D music pose dataset containing 1,408 dance motions tailored to various music genres. FineDance [26], a 3D dataset focusing on fine-grained hand motions, includes 14 hours of dance data. Models like Bailando [38] and EDGE [44] have leveraged AIST++ for training. Bailando [38] pioneered using a VQ-VAE for 3D pose encoding, followed by an Actor-Critic GPT model to generate body poses conditioned on music. EDGE [44] applied a diffusion framework to predict human poses from a noisy sequence, conditioning on music and employing long-form sampling for extended dance generation. DabFusion [55] explores image-to-video generation conditioned on music. Despite these advancements, existing approaches trained on 3D datasets often generate dance poses with limited diversity. X-Dancer overcomes these data constraints by leveraging a broad spectrum of 2D dance motions, aligning complex movements with musical beats, and utilizing widely available video data to enhance scalability.
@@ 2.2 Human Image Animation
With the advancements in diffusion models [16, 14, 5, 52, 48], generating high-quality human videos has become feasible. The introduction of ControlNet [62, 50] and PoseGuider [17, 18] has further empowered prior methods [17, 7, 54, 66, 8] to create realistic dance videos from a reference image using motion signals such as 2D skeletons, dense poses, and depth maps. However, generating expressive, temporally smooth 2D motion sequences that maintain consistency with reference body shapes and align well with music remains an open challenge. Several methods [53, 21, 42] adopt an end-to-end pipeline for generating realistic talking head videos from audio by injecting audio features directly into the network via cross-attention layers. While these methods excel in capturing micro-expressions and lip movements, they fall short in handling highly dynamic and articulated dance movements. X-Dancer implicitly incorporates generated motion tokens into a diffusion backbone to animate the reference image, enabling high-quality and shape-consistent dance video generation.
Figure 2: Overview of X-Dancer. We propose a cross-conditional transformer model to autoregressively generate 2D human poses synchronized with input music, followed by a diffusion model to produce high-fidelity videos from a single reference image $I_{R}.$ First, we develop a multi-part compositional tokenization for 2D poses, encoding and quantizing each body part independently with keypoint confidence. A shared decoder merges these tokens into a whole-body, confidence-aware pose. Next, a GPT-based transformer autoregressively predicts future pose tokens with causal attention, conditioned on past poses and aligned music embeddings, as well as global music styles and prior motion context. A motion decoder is trained to generate multi-scale spatial pose guidance upsampled from a learned feature map, integrating the generated motion tokens within a temporal window (16 frames) via AdaIN. By co-training the motion decoder and temporal modules, our diffusion model is capable of synthesizing temporally smooth and high-fidelity video frames, while maintaining consistent appearance with the reference image with a trained reference net.
@@ 3.1 Data Representations
Video Generation. To generate human dance videos, we employ Latent Diffusion Models [37] that synthesize samples in the image latent space, facilitated by a pretrained image auto-encoder. During training, latent features of images are progressively corrupted with Gaussian noise $\epsilon,$ following the Denoising Diffusion Probabilistic Model (DDPM) framework [15, 39]. A UNet-based denoising framework, enhanced with spatial and temporal attention blocks, is then trained to learn the reverse denoising process.
@@
We apply human-centric cropping to the training videos of half and full-body dances, yielding a unified resolution of $896\times 512$. Rather than modeling intricate pixel-wise movements directly on music, we first establish the correlation between music and human dance movement, which subsequently guides the final visual synthesis.
@@
Motion Modeling. In contrast to prior methods that generate 3D human motions, we represent diverse dance motions as 2D pose sequences. Compared to its 3D counterparts, 2D human dance motions are widely accessible from large collections of monocular videos, eliminating the need for complex multiview capture setups or labor-intensive 3D animations. Furthermore, the 2D pose detection is significantly more reliable and robust. To enhance the realism and expressiveness, we model not only large body articulations but also finer details on head movements and hand gestures. Notably, we incorporate keypoint confidence into our pose representation, allowing the model to account for motion blur and occlusions. Each per-frame whole-body pose with $P$ joints is thus represented as $p\in\mathcal{R}^{P\times 3},$ where the last dimension encodes the keypoint confidence.
@@
Music Embedding. Inspired from [44], we utilize the pre-trained Jukebox model [11] to extract rich musical features, supplemented by rhythmic information with one-hot encoding of music beats using an audio processing toolbox Librosa [22]. We resample and synchronize these extracted embeddings—denoted as $F^{J}_{1:T}$ for Jukebox and $F^{L}_{1:T}$ for Librosa—to the video frame rate, ensuring per-frame alignment between the music and visual elements.
3.2 Transformer-Based Music to Dance Motion
@@
Given a collection of monocular, single-person, music-synchronized dance videos with paired 2D whole-body pose detections, we aim to model the intricate, non-linear correlation between skeletal dance motion and musical features. To achieve this, we first introduce a compositional, confidence-aware VQ-VAE, which captures diverse and nuanced human dance poses across different body parts. Next, we leverage a GPT-like transformer to autoregressively predict future pose tokens, modeling temporal motion token transitions in synchronization with music embeddings.
@@
Compositional Confidence-Aware Human Pose Tokenization. Our approach builds on the standard VQ-VAE framework, trained in a self-supervised manner. Given whole-body 2D poses with associated keypoint confidences $p$, a 1D convolutional encoder maps $p$ into a latent pose embedding $z_{e}(p)=E(p),$ which is then quantized by mapping to its nearest representation $z_{q}(p)$ within a learnable codebook, and finally the decoder $D$ reconstructs the pose $\hat{p}$ from $z_{q}(p).$ However, as prior studies [38, 58, 49] suggest, the dependencies between spatial keypoints are complex, and a vanilla VQ-VAE often struggles to capture subtle pose details, such as finger movements and head tilts, due to information loss during quantization and multi-frequency nature of pose variations across different body parts.
@@
To improve expressive coverage, we train independent 2D pose encoders $E^{j}$ and learn $B=5$ separate codebooks $Z^{j}_{q}$ for upper and lower half bodies, left and right hands, and head respectively, allowing the model to spatially decompose 2D whole-body pose variations across different frequencies. With pose partition, distinct body-part codes can be flexibly combined, enhancing the range of expressiveness represented within each individual codebook. To capture part-wise spatial correlations and ensure information flow across body parts, we concatenate the quantized pose latents and feed them into a shared decoder. The resulting embedding is then mapped to reconstructed keypoint coordinates through separate projection heads, enabling joint reconstruction while preserving nuanced part-specific details.
@@ 3.3 Diffusion-Based Dance Video Synthesis
We employ a diffusion model to synthesize high-resolution, realistic human dance videos, conditioned on the generated motions from our trained transformer together with a given reference image. To achieve this, we leverage a pretrained T2I diffusion backbone [37] and incorporate additional temporal modules [14] for temporal consistency. For transferring the reference image context, a reference network, as a trainable copy of the backbone UNet, extracts reference features of identity appearance and background which are cross-queried by the self-attentions within the backbone UNet. Motion control is achieved through an addition module, often configured in ControlNet [62] or light-weight PoseGuider [17, 18], which translates the motion conditions into 2D spatial guidance additive to the UNet features.
@@
In place of explicit translation of motion tokens into 2D skeleton maps as pose guider conditions, we introduce a trainable motion decoder that implicitly and directly translates the 1D pose tokens into 2D spatial guidance, added towards the intermediate features of the denoising UNet. Starting from a learnable 2D feature map, the decoder injects the 1D pose token sequences including the keypoint confidences with AdaIN layers [19], progressively upsampling this feature map into multiple scales aligned with the resolutions of the denoising UNet features. The motion decoder is trained alongside the temporal module within a 16-frame sequence window, effectively decoding token sequences into continuous pose guidance that integrates temporal context from adjacent frames. Moreover, by incorporating reference image context during training, we observe empirically that the decoded pose guidance retains minimal identity and body shape information compared to explicit 2D poses, enabling generated pose tokens to adapt seamlessly to subjects with varied body shapes and appearances.